# Composition of Scalar Mesons and Their Effects on Nuclear Matter Properties in an Extended Linear Sigma Model

### v2

@@ II.2 Nuclear matter properties and neutron star structures
Using the relations in Eq. (7), the degrees of freedom of the parameter space are reduced to $11$, explicitly, $\alpha$, $\beta$, $c_{4}$, $e_{3}$, $\tilde{h}_{2}$, $\tilde{g}_{3}$, $a_{1}$, $c$, $g$, $h$, and $e$. The physical quantities that can be used to constrain these parameters are the bare masses of the scalar mesons and nucleons, as well as the NM properties around saturation density, $n_{0}$. The empirical constraints and corresponding fitted values are shown in Tabs. 2 and 3, and the fitted parameters are listed in Tab. 1.
Table 1: Values of the parameters used in this work. $\alpha$ and $\beta$ are in units of MeV, and $e_{3}$ is in units of GeV. The other parameters are dimensionless. The parameter sets are denoted as el-g30g, el-g30e, el-g30eg, el-g350eg, el-g3100eg, and el-g3150eg, respectively. ’el’ refers to the bELSM; ’g3’ refers to the four-vector meson coupling in Eq. (3c), with the number following it denoting the corresponding magnitude; ’e’ or ’g’ indicates that the coupling between baryons and scalar mesons is mainly from ’e’ or ’g’ terms in Eq. (3a), and ’eg’ indicates that the contributions of the ’e’ and ’g’ terms are similar.
@@
el-g30g $61.1$ $24.7$ $-2.06$ $44.0$ $80.1$ $1.59$ $-0.792$ $15.4$ $4.25$ $11.4$ $-0.0270$
el-g30e $61.1$ $24.4$ $-2.05$ $43.6$ $80.0$ $0.542$ $11.4$ $0.234$ $4.17$ $-0.790$ $15.1$
el-g30eg $61.4$ $26.4$ $-2.10$ $45.6$ $79.3$ $0.397$ $9.51$ $6.54$ $4.10$ $-2.61$ $8.75$
el-g350eg $61.2$ $25.6$ $-2.10$ $44.4$ $79.9$ $51.5$ $10.1$ $6.35$ $4.14$ $-2.65$ $9.00$
el-g3100eg $60.8$ $24.0$ $-2.09$ $42.4$ $80.8$ $100$ $10.6$ $7.10$ $4.19$ $-2.88$ $8.34$
el-g3150eg $60.7$ $24.3$ $-2.11$ $42.0$ $81.0$ $150$ $11.1$ $7.15$ $4.18$ $-3.05$ $8.30$
Table 2: The spectra of hadrons are in units of $\mathrm{MeV}$. The empirical values are chosen as the real parts of the corresponding resonance T-matrix poles or masses from Ref. Navas et al. (2024). The excited state $f_{0}^{\prime}$ is chosen to be $f_{0}(1370)$ with $a_{0}^{\prime}$ being $a_{0}(1450)$ and $\sigma^{\prime}$ being $f_{0}(1500)$. The constraints of $f_{0}(1370)$ and $a_{0}(1450)$ are combined together, due to $\rm U(3)_{V}$ symmetry. The definition of the parameter set is shown in Tab. 1.
@@
Table 3: Nuclear matter properties at saturation density $n_{0}$. The empirical values and their error bars are set according to analyses from Refs. Sedrakian et al. (2023); Piekarewicz (2004); Colo et al. (2004); Khan and Margueron (2012); Dutra et al. (2012, 2014); Oertel et al. (2017); Tews et al. (2017); Wang et al. (2018); Xie and Li (2019); Choi et al. (2021); Grams et al. (2022). The $n_{0}$ is in units of $\rm fm^{-3}$ and the other properties are in units of $\rm MeV$. The definition of the parameter set is shown in Tab. 1.
@@
Empirical $0.155\pm 0.050$ $-16.0\pm 1.0$ $31.7\pm 3.2$ $-200\pm 600$ $58.7\pm 28.1$ $250\pm 50$
el-g30g $0.155$ $-14.6$ $30.9$ $451$ $83.6$ $418$
el-g30e $0.155$ $-14.6$ $31.6$ $479$ $85.8$ $419$
el-g30eg $0.155$ $-14.6$ $30.1$ $421$ $82.2$ $415$
el-g350eg $0.155$ $-15.2$ $30.9$ $-392$ $80.7$ $370$
el-g3100eg $0.155$ $-15.4$ $31.4$ $-1020$ $71.7$ $317$
el-g3150eg $0.155$ $-15.6$ $31.6$ $-1470$ $63.7$ $253$
It can be seen from Tab. 2 that all the parameter sets reproduce hadron spectra consistent with the results from the Particle Data Group Navas et al. (2024). Regarding the NM properties at $n_{0}$ shown in Tab. 3, the saturation density is consistently fixed at $0.155~{}\rm fm^{-3}$, and $E_{\rm sym}$ aligns with the empirical value of $(30.9\pm 1.9)~{}\rm MeV$. However, the other properties decrease with the magnitude of $\tilde{g}_{3}$. It appears that $\tilde{g}_{3}$ should be approximately $150$ to satisfy the constraints of NM properties.
When we extend our study beyond the saturation density to a broader density range to examine NS structures, these parameter sets result in significant differences. The M-R relations of NSs are shown in Fig. 1, which is computed using the Tolman-Oppenheimer-Volkoff (TOV) equation Tolman (1939); Oppenheimer and Volkoff (1939) with pure hadron EOSs from different models.
@@
As we can see in Fig. 1(a), the model without the $a_{0}$ meson cannot reach the GW170817 constraint due to its overly loose structure of NSs. The models with the $a_{0}$ meson, both el-g30eg and FSU-$\delta 6.7$, can satisfy the GW170817 and MSP J0740+6620 constraints simultaneously, highlighting the importance of the $a_{0}$ meson—an iso-vector scalar meson in NS studies. From Fig. 1(b), it is evident that only EOSs with similar contributions from the $e$ and $g$ terms can meet the observational constraint of MSP J0740+6620; otherwise, the $M_{\rm max}$ will be slightly lower. However, the properties of NM at $n_{0}$ are not sensitive to the contributions of the $e$ and $g$ terms. Regarding the $\tilde{g}_{3}$ terms, as shown in Fig. 1(c), the $M_{\rm max}$ of NSs is sensitive to the magnitude of $\tilde{g}_{3}$: a larger $\tilde{g}_{3}$ coupling—four vector meson couplings—leads to a smaller $M_{\rm max}$ of NSs, even though the NM properties at $n_{0}$ are more consistent with empirical values.
@@
where $N=\binom{p}{n}$ and $\tau_{3}$ is the 3rd component of Pauli matrix. The values of the couplings are list in Table 4.
@@
$g_{\sigma NN}$ $g_{aNN}$ $g_{fNN}$ $g_{\omega NN}$ $g_{\rho NN}$
TM1 Sugahara and Toki (1994) $-10.0$ $-12.6$ $-4.63$
FSU-$\delta$6.7 Li et al. (2022) $10.2$ $6.7$ $-13.4$ $-7.27$
el-g30g $-6.17$ $5.13$ $2.95$ $-6.09$ $5.30$
el-g30e $-6.20$ $-5.03$ $3.00$ $6.09$ $5.30$
el-g30eg $-5.97$ $-0.671$ $2.68$ $6.06$ $3.45$
el-g350eg $-6.12$ $-0.852$ $2.85$ $6.37$ $3.71$
el-g3100eg $-6.36$ $-0.442$ $3.19$ $6.73$ $3.85$
el-g3150eg $-6.38$ $-0.413$ $3.20$ $7.09$ $4.04$
In our bELSM, the magnitudes of the $e$ and $g$ terms determine $g_{aNN}$, while $g_{\sigma NN}$ and $g_{fNN}$ remain almost the same across different parameter sets. However, the M-R relation of NSs is sensitive to $g_{aNN}$: a larger $g_{aNN}$ leads to a smaller $M_{\rm max}$ for NSs, as shown in Fig. 1(b). This is not the case in FSU-$\delta 6.7$, which does not consider $f_{0}$. In bELSM, all possible terms involving the $\delta$ meson at the lowest order are considered, including all three- and four-point interactions: self-couplings, couplings to vector mesons, and couplings to scalar mesons. These terms become increasingly important with density, especially those relate to vector mesons. By including all possible multi-meson interactions at the lowest order, respecting the chiral symmetry pattern at low energies, the $g_{\sigma NN}$ and $g_{\omega NN}$ at densities are slightly smaller than those in Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989); Sugahara and Toki (1994); Li et al. (2022), with $g_{\rho NN}$ remaining almost the same.
@@
As shown in Fig. 2(a), the $E_{\rm sym}(n)$ from el-g30eg and FSU-$\delta$6.7 is soft at intermediate densities, around $2n_{0}$, while stiff at subsaturation densities, around $2/3n_{0}$, to meet the constraints of GW170817 Abbott et al. (2018) and the neutron skin thickness of ${}^{208}\rm Pb$ Adhikari et al. (2021); Reed et al. (2021). However, the $E_{\rm sym}(n)$ from TM1 is almost linearly dependent on the density, which is too stiff at intermediate densities, leading to a too loose NS compared to the observations of GW170817, as shown in Fig. 1(a). This highlights the necessity of including the $a_{0}$ meson in the NS EOS, resulting in a plateau-like structure of $E_{\rm sym}(n)$ at intermediate densities. However, a large $g_{aNN}$ will lead to a decreasing $E_{\rm sym}(n)$ at high densities, as shown in Fig. 2(b), which results in a decrease in the $M_{\rm max}$ of NSs, as mentioned in Fig. 1(b). When comparing the $E_{\rm sym}(n)$ from different $\tilde{g}_{3}$ terms in Fig. 2(c), it seems that $E_{\rm sym}(n)$ is not significantly suppressed by the $\tilde{g}_{3}$ term, but the $M_{\rm max}$ of NSs is even smaller than in the case with a large $g_{aNN}$. The reason lies in the SV behavior at high densities, as presented in Fig. 3.
@@
It can be observed that a large $\tilde{g}_{3}$ makes $v_{s}^{2}$ softer compared to $g_{aNN}$ suppression, as seen in Fig. 3(b) and Fig. 3(c). The soft SV behavior, especially when not approaching the conformal limit throughout the densities, predicts a small $M_{\rm max}$ for NSs, which cannot meet the constraints of MSP J0740+6620.
Now, let’s examine the results in Fig. 1(a):
@@
If we only consider the NM properties and the NS structures, both el-g30eg and FSU-$\delta 6.7$ can meet the phenomenological constraints, with FSU-$\delta 6.7$ appearing more favored. However, bELSM reflects the symmetry pattern of QCD at low energies, while the Walecka-type models lack such consideration but introduce possible couplings artificially. Moreover, considering the SV behavior at high densities, el-g30eg allows the SV to approach the conformal limit. This technically stems from including all possible terms about the related baryons and mesons at the lowest order, but the deeper reason may be its close relation to QCD. All the M-R relations in this work are calculated with parameters determined by the NM properties at $n_{0}$, but the parameter space and the resulting NS structures can be further improved via statistical analysis, including constraints from NS observations, as discussed in our past work Guo et al. (2024a).
@@ II.3 Constituents of the scalar mesons
The above discussions suggest possible magnitudes of the 2-quark and 4-quark constituents of the scalar mesons. We list the percentage of the two-quark component $p_{2}=\left|\cos^{2}\theta\right|$ and four-quark component $p_{4}=\left|\sin^{2}\theta\right|$ defined by Eq. (8) in Tab. 5.
Table 5: The possible percentage of the 2-quark and 4-quark constituents of the scalar mesons.
$\sigma$ $\delta$
@@
It is found that the $\sigma$ meson, a flavor singlet, is composed of 2-quark and 4-quark constituents in similar proportions, while the $\delta$ meson, a flavor octet, is primarily composed of 4-quark constituents. This conclusion is drawn after determining the bare masses of the relevant hadrons in Tab. 2 and the NM properties in Tab. 3. This finding is consistent with previous analyses in vacuum Fariborz et al. (2008). When increasing the magnitude of the $\tilde{g}_{3}$ term, the 4-quark constituent percentages of both the $\sigma$ and $\delta$ mesons decrease to find an optimal parameter set that meets empirical constraints. Conversely, when the contributions of the $e$ and $g$ terms are made similar, the 4-quark constituent percentages of both the $\sigma$ and $\delta$ mesons increase.
Regarding the parameter space found in our previous work Ma and Ma (2024), the current optimal parameter spaces are more favored, not only for their consistent NM properties at $n_{0}$ and reasonable EOS behaviors over a wider density region, as discussed above, but also for their prediction of the pion decay constant $f_{\pi}$. The pion decay constant can be obtained using the PCAC relation in the chiral limit (no explicit symmetry breaking due to quark mass) Fariborz et al. (2008): $f_{\pi}=2\sqrt{\alpha^{2}+\beta^{2}}$. The el-g30eg parameter set indicates that $f_{\pi}\approx 134~{}\rm MeV$, which is consistent with recent lattice QCD results Aoki et al. (2020), while $f_{\pi}\sim(60-70)~{}\rm MeV$ was obtained from the parameter space in our previous work. This again highlights the importance of the new flavor structures in the phenomenological analysis to achieve a realistic parameter space.
@@
$\mathcal{L}_{\rm RMF}$ $=$ $C_{\sigma\delta^{2}}\sigma a_{0}^{2}+C_{\sigma^{2}\delta^{2}}\sigma^{2}a_{0}^{2}.$ (10)
Then, according to Tab. 1 and Lagrangian (3b), the corresponding values of the couplings are listed in Tab. 6 for comparison.
Table 6: The values of $C_{\sigma\delta^{2}}$ and $C_{\sigma^{2}\delta^{2}}$ in different models.
$C_{\sigma\delta^{2}}$(MeV) $C_{\sigma^{2}\delta^{2}}$
Zabari-19 Zabari et al. (2019) ±$\pm 1.77$$\pm 0.004$
@@
The larger magnitude of the couplings in Tab. 6 indicates that these multi-meson interactions contribute more to the EOS at both saturation density and high densities, and links the behavior of both regions more closely. This is the reason why we must increase the $K(n_{0})$ to achieve a reasonable NS structures, as shown Fig. 1(c). While the parameter spaces of FSU-$\delta 6.7$ and Zabari-19 make NM properties at $n_{0}$ do not change obviously after introducing these scalar couplings, they affect the EOS at high densities much. This leads to the fact that they can handle the NM properties at $n_{0}$ and NS structures separately, compared to the bELSM. However, since the scalar meson couplings in $\mathcal{L}_{\rm M}$ are determined by $\alpha$, $\beta$, $e_{3}$, and $c_{4}$, which are closely related to the quark constituents of scalar mesons listed in Tab. 5 and important physical values, such as $f_{\pi}$. In contrast, the Walecka-type models, as we are addressing, only introduce possible freedoms and choose operators artificially, which cannot be directly linked to QCD. Therefore, the results in Tab. 6 highlight the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analysis and including operators systematically, especially for phenomenological studies at high densities.
III Conclusion and discussion

### v3

@@ II.2 Nuclear matter properties and neutron star structures
Using the relations in Eq. (7), the degrees of freedom of the parameter space are reduced to $11$, explicitly, $\alpha$, $\beta$, $c_{4}$, $e_{3}$, $\tilde{h}_{2}$, $\tilde{g}_{3}$, $a_{1}$, $c$, $g$, $h$, and $e$. The physical quantities that can be used to constrain these parameters are the bare masses of the scalar mesons and nucleons, as well as the NM properties around saturation density, $n_{0}$. The empirical constraints and corresponding fitted values are shown in Tables 2 and 3, and the fitted parameters are listed in Table 1.
Table 1: Values of the parameters used in this work. $\alpha$ and $\beta$ are in units of MeV, and $e_{3}$ is in units of GeV. The other parameters are dimensionless. The parameter sets are denoted as el-g30g, el-g30e, el-g30eg, el-g350eg, el-g3100eg, and el-g3150eg, respectively; ”el” refers to the bELSM; ”g3” refers to the four-vector meson coupling in Eq. (3c), with the number following it denoting the corresponding magnitude; ”e” or ”g” indicates that the coupling between baryons and scalar mesons is mainly from ”e” or ”g” terms in Eq. (3a), and ”eg” indicates that the contributions of the ”e” and ”g” terms are similar.
@@
el-g30g $61.1$ $24.7$ 2.06-2.06 $44.0$ $80.1$ $1.59$ 0.792-0.792 $15.4$ $4.25$ $11.4$ 0.0270-0.0270
el-g30e $61.1$ $24.4$ 2.05-2.05 $43.6$ $80.0$ $0.542$ $11.4$ $0.234$ $4.17$ 0.790-0.790 $15.1$
el-g30eg $61.4$ $26.4$ 2.10-2.10 $45.6$ $79.3$ $0.397$ $9.51$ $6.54$ $4.10$ 2.61-2.61 $8.75$
el-g350eg $61.2$ $25.6$ 2.10-2.10 $44.4$ $79.9$ $51.5$ $10.1$ $6.35$ $4.14$ 2.65-2.65 $9.00$
el-g3100eg $60.8$ $24.0$ 2.09-2.09 $42.4$ $80.8$ $100$ $10.6$ $7.10$ $4.19$ 2.88-2.88 $8.34$
el-g3150eg $60.7$ $24.3$ 2.11-2.11 $42.0$ $81.0$ $150$ $11.1$ $7.15$ $4.18$ 3.05-3.05 $8.30$
Table 2: The spectra of hadrons are in units of $\mathrm{MeV}$. The empirical values are chosen as the real parts of the corresponding resonance T-matrix poles or masses from Ref. Navas et al. (2024). The excited state $f_{0}^{\prime}$ is chosen to be $f_{0}(1370)$ with $a_{0}^{\prime}$ being $a_{0}(1450)$ and $\sigma^{\prime}$ being $f_{0}(1500)$. The constraints of $f_{0}(1370)$ and $a_{0}(1450)$ are combined together, due to $\rm U(3)_{V}$ symmetry. The definition of the parameter set is shown in Table 1.
@@
Table 3: Nuclear matter properties at saturation density $n_{0}$. The empirical values and their error bars are set according to analyses from Refs. Sedrakian et al. (2023); Piekarewicz (2004); Colo et al. (2004); Khan and Margueron (2012); Dutra et al. (2012, 2014); Oertel et al. (2017); Tews et al. (2017); Wang et al. (2018); Xie and Li (2019); Choi et al. (2021); Grams et al. (2022). The $n_{0}$ is in units of $\rm fm^{-3}$ and the other properties are in units of $\rm MeV$. The definition of the parameter set is shown in Table 1.
@@
Empirical $0.155\pm 0.050$ $16.0\pm 1.0$ $31.7\pm 3.2$ $200\pm 600$ $58.7\pm 28.1$ $250\pm 50$
el-g30g $0.155$ 14.6-14.6 $30.9$ $451$ $83.6$ $418$
el-g30e $0.155$ 14.6-14.6 $31.6$ $479$ $85.8$ $419$
el-g30eg $0.155$ 14.6-14.6 $30.1$ $421$ $82.2$ $415$
el-g350eg $0.155$ 15.2-15.2 $30.9$ 392-392 $80.7$ $370$
el-g3100eg $0.155$ 15.4-15.4 $31.4$ 1020-1020 $71.7$ $317$
el-g3150eg $0.155$ 15.6-15.6 $31.6$ 1470-1470 $63.7$ $253$
It can be seen from Table 2 that all the parameter sets reproduce hadron spectra consistent with the results from the Particle Data Group Navas et al. (2024). Regarding the NM properties at $n_{0}$ shown in Table 3, the saturation density is consistently fixed at $0.155~\rm fm^{-3}$, and $E_{\rm sym}$ aligns with the empirical value of $(30.9\pm 1.9)~\rm MeV$. However, the other properties decrease with the magnitude of $\tilde{g}_{3}$. It appears that $\tilde{g}_{3}$ should be approximately $150$ to satisfy the constraints of NM properties.
When we extend our study beyond the saturation density to a broader density range to examine NS structures, these parameter sets result in significant differences. The M-R relations of NSs are shown in Fig. 1, which is computed using the Tolman-Oppenheimer-Volkoff equation Tolman (1939); Oppenheimer and Volkoff (1939) with pure hadron EOS from different models.
@@
As we can see in Fig. 1(a), the model without the $a_{0}$ meson cannot reach the GW170817 constraint due to its overly loose structure of NSs. The models with the $a_{0}$ meson, both el-g30eg and FSU-$\delta 6.7$, can satisfy the GW170817 and MSP J0740+6620 constraints simultaneously, highlighting the importance of the $a_{0}$ meson—an isovector scalar meson in NS studies. From Fig. 1(b), it is evident that only EOS with similar contributions from the $e$ and $g$ terms can meet the observational constraint of MSP J0740+6620; otherwise, the $M_{\rm max}$ will be slightly lower. However, the properties of NM at $n_{0}$ are not sensitive to the contributions of the $e$ and $g$ terms. Regarding the $\tilde{g}_{3}$ terms, as shown in Fig. 1(c), the $M_{\rm max}$ of NSs is sensitive to the magnitude of $\tilde{g}_{3}$: a larger $\tilde{g}_{3}$ coupling—four vector meson couplings—leads to a smaller $M_{\rm max}$ of NSs, even though the NM properties at $n_{0}$ are more consistent with empirical values.
@@
where $N=\binom{p}{n}$ and $\tau_{3}$ is the third component of Pauli matrix. The values of the couplings are list in Table 4.
@@
$g_{\sigma NN}$ $g_{{a_{0}NN}}$ $g_{fNN}$ $g_{\omega NN}$ $g_{\rho NN}$
TM1 Sugahara and Toki (1994) 10.0-10.0 12.6-12.6 4.63-4.63
FSU-$\delta$6.7 Li et al. (2022) $10.2$ $6.7$ 13.4-13.4 7.27-7.27
el-g30g 6.17-6.17 $5.13$ $2.95$ 6.09-6.09 $5.30$
el-g30e 6.20-6.20 5.03-5.03 $3.00$ $6.09$ $5.30$
el-g30eg 5.97-5.97 0.671-0.671 $2.68$ $6.06$ $3.45$
el-g350eg 6.12-6.12 0.852-0.852 $2.85$ $6.37$ $3.71$
el-g3100eg 6.36-6.36 0.442-0.442 $3.19$ $6.73$ $3.85$
el-g3150eg 6.38-6.38 0.413-0.413 $3.20$ $7.09$ $4.04$
In our bELSM, the magnitudes of the $e$ and $g$ terms determine $g_{aNN}$, while $g_{\sigma NN}$ and $g_{fNN}$ remain almost the same across different parameter sets. However, the M-R relation of NSs is sensitive to $g_{aNN}$: a larger $g_{aNN}$ leads to a smaller $M_{\rm max}$ for NSs, as shown in Fig. 1(b). This is not the case in FSU-$\delta 6.7$, which does not consider $f_{0}$. In bELSM, all possible terms involving the $a_{0}$} meson at the lowest-order are considered, including all three- and four-point interactions: self-couplings, couplings to vector mesons, and couplings to scalar mesons. These terms become increasingly important with density, especially those relate to vector mesons. By including all possible multimeson interactions at the lowest-order, respecting the chiral symmetry pattern at low energies, the $g_{\sigma NN}$ and $g_{\omega NN}$ at densities are slightly smaller than those in Walecka-type models Holinde and Machleidt (1975); Erkelenz (1974); Nagels et al. (1978); Machleidt (1989); Sugahara and Toki (1994); Li et al. (2022), with $g_{\rho NN}$ remaining almost the same.
@@
As shown in Fig. 2(a), the $E_{\rm sym}(n)$ from el-g30eg and FSU-$\delta$6.7 is soft at intermediate densities, around $2n_{0}$, while stiff at subsaturation densities, around $2/3n_{0}$, to meet the constraints of GW170817 Abbott et al. (2018) and the neutron skin thickness of $^{208}\rm Pb$ Adhikari et al. (2021); Reed et al. (2021). However, the $E_{\rm sym}(n)$ from TM1 is almost linearly dependent on the density, which is too stiff at intermediate densities, leading to too-loose NS compared to the observations of GW170817, as shown in Fig. 1(a). This highlights the necessity of including the $a_{0}$ meson in the NS EOS, resulting in a plateaulike structure of $E_{\rm sym}(n)$ at intermediate densities. However, a large $g_{aNN}$ will lead to a decreasing $E_{\rm sym}(n)$ at high densities, as shown in Fig. 2(b), which results in a decrease in the $M_{\rm max}$ of NSs, as mentioned in Fig. 1(b). When comparing the $E_{\rm sym}(n)$ from different $\tilde{g}_{3}$ terms in Fig. 2(c), it seems that $E_{\rm sym}(n)$ is not significantly suppressed by the $\tilde{g}_{3}$ term, but the $M_{\rm max}$ of NSs is even smaller than in the case with a large $g_{aNN}$. The reason lies in the SV behavior at high densities, as presented in Fig. 3.
@@
It can be observed that a large $\tilde{g}_{3}$ makes $v_{s}^{2}$ softer compared to $g_{{a_{0}NN}}$ suppression, as seen in Figs. 3(b) and 3(c). The soft SV behavior, especially when not approaching the conformal limit throughout the densities, predicts a small $M_{\rm max}$ for NSs, which cannot meet the constraints of MSP J0740+6620.
Now, let us examine the results in Fig. 1(a):
@@
If we only consider the NM properties and the NS structures, both el-g30eg and FSU-$\delta 6.7$ can meet the phenomenological constraints, with FSU-$\delta 6.7$ appearing more favored. However, bELSM reflects the symmetry pattern of QCD at low energies, while the Walecka-type models lack such consideration but introduce possible couplings artificially. Moreover, considering the SV behavior at high densities, el-g30eg allows the SV to approach the conformal limit. This technically stems from including all possible terms about the related baryons and mesons at the lowest-order, but the deeper reason may be its close relation to QCD. All the M-R relations in this work are calculated with parameters determined by the NM properties at $n_{0}$, but the parameter space and the resulting NS structures can be further improved via statistical analysis, including constraints from NS observations, as discussed in our past work Guo et al. (2024a).
@@ II.3 Constituents of the scalar mesons
The above discussions suggest possible magnitudes of the two-quark and four-quark constituents of the scalar mesons. We list the percentage of the two-quark component $p_{2}=\left|\cos^{2}\theta\right|$ and four-quark component $p_{4}=\left|\sin^{2}\theta\right|$ defined by Eq. (8) in Table 5.
Table 5: The possible percentage of the two-quark and four-quark constituents of the scalar mesons.
$\sigma$ $a_{0}$}
@@
It is found that the $\sigma$ meson, a flavor singlet, is composed of two-quark and four-quark constituents in similar proportions, while the $a_{0}$} meson, a flavor octet, is primarily composed of four-quark constituents. This conclusion is drawn after determining the bare masses of the relevant hadrons in Table 2 and the NM properties in Table 3. This finding is consistent with previous analyses in vacuum Fariborz et al. (2008). When increasing the magnitude of the $\tilde{g}_{3}$ term, the four-quark constituent percentages of both the $\sigma$ and $a_{0}$} mesons decrease to find an optimal parameter set that meets empirical constraints. Conversely, when the contributions of the $e$ and $g$ terms are made similar, the four-quark constituent percentages of both the $\sigma$ and $a_{0}$} mesons increase.
Regarding the parameter space found in our previous work Ma and Ma (2024), the current optimal parameter spaces are more favored, not only for their consistent NM properties at $n_{0}$ and reasonable EOS behaviors over a wider density region, as discussed above but also for their prediction of the pion decay constant $f_{\pi}$. The pion decay constant can be obtained using the partially conserved axial current (PCAC) relation in the chiral limit (no explicit symmetry breaking due to quark mass) Fariborz et al. (2008): $f_{\pi}=2\sqrt{\alpha^{2}+\beta^{2}}$. The el-g30eg parameter set indicates that $f_{\pi}\approx 134~\rm MeV$, which is consistent with recent lattice QCD results Aoki et al. (2020), while $f_{\pi}\sim(60-70)~\rm MeV$ was obtained from the parameter space in our previous work. This again highlights the importance of the new flavor structures in the phenomenological analysis to achieve a realistic parameter space.
@@
$\mathcal{L}_{\rm RMF}$ $=$ $C_{\sigma{a_{0}}^{2}}\sigma a_{0}^{2}+C_{\sigma^{2}{a_{0}}^{2}}\sigma^{2}a_{0}^{2}.$ (10)
Then, according to Table 1 and Lagrangian (3b), the corresponding values of the couplings are listed in Table 6 for comparison.
Table 6: The values of $C_{\sigma{a_{0}}^{2}}$ and $C_{\sigma^{2}{a_{0}}^{2}}$ in different models.
$C_{\sigma{a_{0}}^{2}}$(MeV) $C_{\sigma^{2}{a_{0}}^{2}}$
Zabari-19 Zabari et al. (2019) $\pm 1.77$ $\pm 0.004$
@@
The larger magnitude of the couplings in Table 6 indicates that these multimeson interactions contribute more to the EOS at both saturation density and high densities, and links the behavior of both regions more closely. This is the reason why we must increase the $K(n_{0})$ to achieve a reasonable NS structures, as shown Fig. 1(c). While the parameters from FSU-$\delta$6.7 and Zabari-19 make NM properties at $n_{0}$ do not change obviously compared to those without couplings between the $a_{0}$ and $\sigma$ mesons, for the small magnitudes of these couplings, they affect the EOS at high densities much. This leads to the fact that they can handle the NM properties at $n_{0}$ and NS structures separately, compared to the bELSM. However, since the scalar meson couplings in $\mathcal{L}_{\rm M}$ are determined by $\alpha$, $\beta$, $e_{3}$, and $c_{4}$, which are closely related to the quark constituents of scalar mesons listed in Table 5 and important physical values, such as $f_{\pi}$. In contrast, the Walecka-type models, as we are addressing, only introduce possible freedoms and choose operators artificially, which cannot be directly linked to QCD. Therefore, the results in Table 6 highlight the importance of considering the chiral symmetry pattern of QCD at low energies in phenomenological analysis and including operators systematically, especially for phenomenological studies at high densities.
III Conclusion